# Target-driven Self-Distillation for Partial Observed Trajectories Forecasting

### v1

@@ 4.5 Training and Inference
Table 3: Comparison of the minADE6/minFDE6/MR6 performance of different ablation versions of TSD-H. The best results are highlighted in bold.
@@
Table 4: Comparison of the minADE6/minFDE6/MR6 performance of different ablation versions of self-distillation part of TSD-H. The best results are highlighted in bold.
5 Experiments
5.1 Experiment Settings
Dataset. We evaluated our proposed model on the Argoverse1 Chang et al. (2019) and Argoverse2 Wilson et al. (2021) motion forecasting datasets. The Argoverse1 dataset provides agent trajectories and high-definition map data, containing 323,557 real driving scenarios divided into training, validation, and test sets with 205,942, 39,472, and 78,143 samples, respectively. All training and validation scenes consist of 5-second sequences sampled at 10 Hz, while the test set only reveals the first 2 seconds of the trajectories. Based on the initial 2 seconds of observation, the model is required to predict the agent’s movement over the next 3 seconds. The Argoverse2 dataset provides agent trajectory data sampled at 10Hz for up to 11 seconds. The training, validation, and test sets contain 199,908, 24,988, and 24,984 samples. Each sample includes the trajectory of the agent to be predicted, and the trajectories of surrounding agents, and comes with high-definition map information of the scene. The objective of the Argoverse2 motion forecasting challenge is to predict the interested agent’s future trajectory for the next 6 seconds based on its trajectory in the previous 5 seconds.
Metrics. We adopted evaluation metrics commonly used in previously state-of-the-art methods Zhou et al. (2023); Wang et al. (2023b, a); Feng et al. (2023), i.e., minimum Average Displacement Error (minADEK), minimum Final Displacement Error (minFDEK), Brier-minimum Final Displacement Error (b-minFDEK), and Miss Rate (MRK). The minADE evaluates the $L2$ distance between the coordinates of each step of the top K best-predicted future trajectories by the model and the ground truth, averaged over all time steps. The minFDE evaluates the accuracy of the last point’s predicted coordinates. To better gauge the accuracy of uncertainty predictions, b-minFDE introduces a weighting factor $(1-\hat{\pi})^{2}$ to FDE, where $\hat{\pi}$ represents the probability score allocated by the model to the best-predicted result. The MR metric evaluates the percentage of predictions in minFDE that exceed 2 meters. Typically, K is set to 1 and 6. Only the top K predicted trajectories are evaluated if the model outputs more than K trajectories.
Implement Details. We trained HiVT and QCNet integrated with our method (hereinafter referred to as HiVT and QCNet) using the PyTorch deep learning library Paszke et al. (2019) on 8 RTX 3090 GPUs. The training process included 64 epochs and utilized the AdamW optimizer Loshchilov and Hutter (2018). The batch size for HiVT was set to 32, while it was set to 16 for QCNet due to memory constraints. The initial learning rate, weight decay, and dropout rate for HiVT were set to 3e-4, 1e-4, and 0.1, respectively; for QCNet, the initial learning rate was set to 5e-4, the weight decay to 1e-4, and the dropout rate to 0.1. Both models were trained using a cosine annealing scheduler Loshchilov and Hutter (2016) to adjust the learning rate. The architecture of HiVT consists of 1 layer of agent-agent and agent-lane interaction modules, 4 layers of temporal learning modules, and 3 layers of global interaction modules. For QCNet, the number of cross-attention and self-attention layers was set to 2, the MLP contains 3 layers, and the hidden dimension size is set to 128.
5.2 Quantitative Comparisons
Figure 3: Qualitative results of HiVT and TSD-H. The past trajectories are shown in yellow, the ground-truth trajectories are shown in red, and the predicted trajectories are shown in green. The white boxes are other vehicles around, and the red boxes are agents
Overall Performance with State-of-the-Art. We compare the proposed model with state-of-the-art motion forecasting models on the Argoverse2 motion forecasting dataset Wilson et al. (2021). As shown in Table 1, the results indicate that the proposed model exhibits superior performance compared to state-of-the-art methods such as GANet Wang et al. (2023a), QCNet Zhou et al. (2023), MacFormer Feng et al. (2023), Gnet Gao et al. (2023), ProphNet Wang et al. (2023b), and ForecastMAE Cheng et al. (2023). In the ensemble version of TSD-Q, we employed a simple bagging Breiman (1996) ensemble learning technique to enhance the performance of our model by aggregating the predictions from 5 variants of our proposed model (with different numbers of layers, prediction ranges, and random seeds), resulting in trajectories far exceeding the required prediction output. Subsequently, we applied the weighted K-Means Huang et al. (2005) to cluster these trajectories into the same number of categories as stipulated by the Argoverse 2 motion forecasting challenge (with $K=6$), adopting the predicted $\pi$ as the weight. We then calculated the mean of the trajectories within each category to represent the final trajectory of each category.
Robustness with State-of-the-Art. On the Argoverse 1 dataset , we compared the robustness of state-of-the-art models under partial observation trajectory inputs. As shown in Figure 1, we built TSD-H based on HiVT and compared it with HiVT and the state-of-the-art distillation-based method, POP, under different hidden dimension settings. The results indicate that when the input is partial observation data, the performance of HiVT significantly deteriorates regardless of whether the hidden dimension is 64 or 128. Although POP-H can improve model performance under partial observations, it degrades performance under full observations. In contrast, our proposed TSD-H not only significantly enhances the model’s robustness under partial observations at hidden dimensions of 64 or 128 but also maintains or even improves performance under full observations.
5.3 Ablation Studies
@@
Component Ablation. In Table 3, we analyze the gains of each component of the proposed TSD based on TSD-H. The contents of the table show that by predicting accurate proxy target points and performing target-guided trajectory prediction, the model achieves improvements in both fully observed and partially observed trajectories, resulting in an overall performance enhancement. After adding self-distillation to the motion forecasting training, the model shows further improvement in fully observed trajectories and a significant increase in robustness for partially observed trajectory inputs.
@@
Target Number. We investigate the impact of the number of generated target points on the performance of the proposed model, comparing the results of predictions with six target points, three target points, and a single target point. When testing variants with more than three target points, the number of iterations in the trajectory prediction module is also increased to match the number of target points. As shown in Table 5, when only generating a single long-term target point to guide trajectory generation, the broad time horizon leads to inaccurate target points, significantly impairing the trajectory prediction. When too many target points are predicted, the error accumulation results in less accurate target point generation, slightly reducing trajectory prediction performance.
5.4 Qualitative Results
To demonstrate the effectiveness of our method, we conducted qualitative experiments on the Argoverse1 dataset, as shown in 3. Case 1 shows a straight-driving scenario with an input history length of 20. Case 2 shows a turning scenario with an input history length of 20. It can be observed that our method, TSD, outperforms the original HIVT in both scenarios. In Case 1, the predicted trajectory of our method is closer to the ground truth than HIVT’s prediction. In Case 2, HIVT incorrectly predicts a straight path, while our method successfully predicts the turning scenario. Case 3 shows an experiment where only part of the historical trajectory is provided, with a history length of 10. In this case, our method predicts the turning trajectory more closely to the ground truth compared to the original HIVT.
Targets Iterations $\mathrm{minADE}_{6}\downarrow$ $\mathrm{minFDE}_{6}\downarrow$ $\mathrm{MR_{6}}\downarrow$
1 1 0.74 1.33 0.17
1 3 0.74 1.33 0.17
1 6 0.74 1.32 0.17
3 3 0.70 1.24 0.15
3 6 0.70 1.24 0.16
6 6 0.71 1.26 0.16
Table 5: Comparison of the impact of different numbers of target points and the number of iterations for predicting trajectories on the final performance of the proposed model. We investigated as many combinations as possible to prove that our chosen target points and the number of rounds for trajectory prediction are optimal.
@@ 6 Conclution
This paper presented a self-distillation motion forecasting method based on anchor-free target point guidance, tailored to the characteristics of partial observation trajectory prediction tasks. First, the method predicted accurate long-term and short-term target points for different time horizons, which guided precise trajectory prediction under partial observation conditions. Second, we designed a self-distillation approach for partial observation trajectories, ensuring that the feature distributions and prediction results are consistent between full and partial observation scenarios. Extensive experiments demonstrate that the trained model can achieve more accurate motion forecastings under partial observation and also improve prediction performance under full observation.
A Target Generator
Figure 4 illustrates the internal structure of the proposed sequential target generator. This generator is based on the Transformer Vaswani et al. (2017) decoder architecture, utilizing cross-attention mechanisms to enable interaction between the query and both road features and agent trajectory features. The final predictions for long-term and short-term target points are produced through feed-forward layer.
Figure 4: Sequential target generator.
B Loss Computation
BT $\mathcal{L}_{\text{cls}}$ with T-NLL $\mathrm{minADE}_{6}\downarrow$ $\mathrm{minFDE}_{6}\downarrow$ $\mathrm{MR_{6}}\downarrow$
0.73 1.31 0.17
$\checkmark$ 0.72 1.28 0.16
$\checkmark$ 0.71 1.25 0.16
$\checkmark$ $\checkmark$ 0.70 1.24 0.15
Table 6: Comparison of the impact of different gradient backpropagation strategies and training losses on the performance of the proposed model. BT denote choose the mode with best target prediction as the best mode. $\mathcal{L}_{\text{cls}}$ with T-NLL refers to calculating $\mathcal{L}_{\text{cls}}$ based on the negative log-likelihood loss of trajectory.
Ablation Obs. = 1 Obs. = 5 Obs. = 10 Obs. = 15 Obs. = 20
w/ KL Loss 1.047/1.615/0.216 0.740/1.110/0.116 0.697/1.044/0.103 0.682/1.014/0.098 0.678/1.008/0.099
w/o KL Loss 1.040/1.600/0.211 0.739/1.098/0.117 0.694/1.031/0.104 0.679/1.009/0.098 0.675/0.998/0.096
Table 7: Comparison of the minADE6/minFDE6/MR6 performance of TSD-H self-distillation with and without the addition of KL Loss. The best results are highlighted in bold.
We investigate the impact of different gradient backpropagation strategies and different loss functions for the mixture factor $\pi$ on the prediction accuracy of target points and the overall performance of motion forecasting, respectively. In the winner-takes-all training strategy, we choose the mode with the best target point prediction for gradient backpropagation and explore the impact of selecting the mode with the best trajectory prediction. For the mixture factor $\pi$, as per $\mathcal{L}_{cls}$, its objective function is constructed based on the negative log-likelihood loss of the trajectory. To explore the impact of different objective functions, we test constructing $\mathcal{L}_{\text{cls}}$ based on the negative log-likelihood loss of the target points.
As shown in Table 6, selecting the mode with the best prediction of target points for gradient backpropagation significantly enhances the accuracy of target point prediction. This allows the model to generate target points more accurately with the corresponding trajectories. Moreover, using the loss function $\mathcal{L}_{cls}$ also improves the overall performance of the proposed model without significantly affecting the prediction accuracy of target points.
C External Analysis
Effect of KL Divergence. Since the self-distillation for classification tasks in Xu and Liu (2019) includes KL loss on the predicted classification results, we also attempted to introduce KL loss between our mode classification results. The results are shown in Table 1. It can be seen that after adding the KL loss, there is a slight decrease in the model’s robustness and overall performance, with improvements in the MR6 metric only for certain observation lengths. We speculate that this is because the addition of KL loss causes the model to focus more on the consistency of the best mode selection under fully observed and partially observed conditions, while reducing its attention to other optimization objectives. In the future, we will explore how to balance various losses and how to compute the loss between fully observed and partially observed predictions more effectively to ensure their consistency.
D Discussion
In the future, we will further explore the self-distillation partial observation motion forecasting method with target point guidance. Regarding target point guidance, we believe that introducing statistical priors of future trajectories can enhance the accuracy of target points when observations are insufficient. In terms of self-distillation, we will investigate how to unify the distributions of predicted trajectories and probabilities between full and partial observation conditions. We hope that these further explorations will enhance the robustness of the model under partial observation and contribute to safer autonomous driving technology.

### v2

@@ 5.1 Experiment Settings
Table 3: Comparison of the minADE6/minFDE6/MR6 performance of different ablation versions of HiVT-TSD. The best results are highlighted in bold.
@@
Table 4: Comparison of the minADE6/minFDE6/MR6 performance of different ablation versions of self-distillation part of HiVT-TSD. The best results are highlighted in bold.
5.2 Result and Quantitative Comparisons
Robustness with Random Masking: We systematically evaluated the motion prediction performance of the proposed method across multiple datasets. As shown in Table 1, we developed three enhanced versions—HiVT-TSD, DeMo-TSD, and LAformer-TSD—based on the mainstream model frameworks HiVT, DeMo, and LAformer, respectively, and conducted comparative experiments with the state-of-the-art knowledge distillation-based method, POP.In addition, we include Forecast-mae in our comparison as a representative mask-based self-supervised pretraining method.
The experimental results demonstrate that as the mask ratio increases under partially observed data with random masking, the prediction performance of all models significantly declines. Notably, while the POP method yields some performance improvements over the original models under random masking conditions, it comes at the cost of reduced prediction accuracy under full observation. Moreover, in certain model architectures, the performance under random masking is even worse than that of the original models (LAformer-POP at masking rates of 60%–80%). In contrast, our proposed TSD method not only significantly outperforms the POP method under high mask ratios but also fully preserves or even slightly enhances the original model performance under complete observation data.On the Argoverse 2 dataset, the results of Forecast-mae show that its performance drops sharply as the masking ratio increases. This indicates that although Forecast-mae benefits from learning structured masking-trajectory reconstruction during pretraining, the learned priors fail to generalize well to the random masking patterns encountered at inference time, resulting in limited robustness under severe observation loss.
To further verify the reliability of these improvements, we conducted paired t-tests across five masking ratios for each model. The results show that the improvements of HiVT (p = 0.0241) and DeMo (p = 0.0336) are statistically significant, while LAformer exhibits a similar trend (p = 0.0656 $\approx$ 0.05). These analyses confirm that the observed gains are consistent and not due to random fluctuations.
Overall, these findings validate the advantages of the TSD method: (1) strong robustness across different levels of observation missingness and consistently outperforms both POP and Forecast-mae under all masking ratios; (2) consistent and statistically reliable performance improvement without sacrificing the base model’s accuracy; and (3) broad adaptability across different motion prediction architectures.
Robustness with Continuous Masking. We conducted a comparative experiment on the robustness of the HiVT, HiVT-POP, and HiVT-TSD models under continuous masking conditions on the Argoverse 1 dataset, as shown in Table 2. The experimental results indicate that, in partially observed scenarios, as the effective observation length decreases (from obs=20 to obs=1), the prediction performance of the original HiVT model exhibits a significant decline, a trend consistent across different hidden dimension settings (64 and 128). Although HiVT-POP shows some improvement over the baseline model at medium observation lengths (Obs= 5 to obs = 15), the enhancement is limited, and its performance under extremely short observation conditions (Obs=1) is even inferior to the baseline model, revealing the method’s limitations in extreme scenarios. In contrast, our proposed TSD method performs excellently across all observation lengths, particularly outperforming HiVT-POP significantly under extremely short input lengths. These results fully validate the robustness of our method against continuous observation missingness and its adaptability to extremely short input lengths.
@@ 5.2 Result and Quantitative Comparisons
Component Ablation. In Table 3, we analyze the gains of each component of the proposed TSD based on HiVT-TSD. The contents of the table show that by predicting accurate proxy target points and performing target-guided trajectory prediction, the model achieves improvements in both fully observed and partially observed trajectories, resulting in an overall performance enhancement. After adding self-distillation to the motion forecasting training, the model shows further improvement in fully observed trajectories and a significant increase in robustness for partially observed trajectory inputs.
@@
Qualitative Result. To demonstrate the effectiveness of our method, we conducted qualitative experiments on the Argoverse1 dataset, as shown in 3. Case 1 shows a straight-driving scenario with an input history length of 20. Case 2 shows a turning scenario with an input history length of 20. It can be observed that our method, TSD, outperforms the original HIVT in both scenarios. In Case 1, the predicted trajectory of our method is closer to the ground truth than HIVT’s prediction. In Case 2, HIVT incorrectly predicts a straight path, while our method successfully predicts the turning scenario. Case 3 shows an experiment where only part of the historical trajectory is provided, with a history length of 10. In this case, our method predicts the turning trajectory more closely to the ground truth compared to the original HIVT.
Model Method Training time Memory Parameter
HiVT origin 32h 6G 0.65M
HiVT POP 64h 8G 1.3M
HiVT TSD 35h 6G 1.3M
DeMo origin 9h 12G 5.9M
DeMo POP 20h 18G 11.8M
DeMo TSD 12h 12G 6M
Table 5: Computational cost compared to other methods.
Computational cost compared to other method. We compared the computational cost of our method on different models in Table 5.Experiments on an NVIDIA RTX 4090 show that TSD provides a more favorable trade-off between efficiency and model complexity. Compared with baseline and POP variants, TSD consistently lowers both training time and memory usage while keeping parameter counts comparable. On DeMo, TSD reduces training time by 40% and memory usage by 33%, with similar improvements observed on HiVT. These results demonstrate that TSD offers a more efficient optimization strategy, achieving faster convergence and lower resource consumption without compromising model capacity.
@@ 6 Conclution
This paper presents a self-distillation motion forecasting method based on anchor-free target point guidance, tailored to the characteristics of partial observation trajectory prediction tasks. First, the method predicts long-term and short-term target points for different time horizons, which guide precise trajectory prediction under partial observation conditions. Second, we design a self-distillation approach for partial observation trajectories, ensuring that the feature distributions and prediction results remain consistent between full and partial observation scenarios. Extensive experiments demonstrate that the trained model achieves more accurate motion forecasting under partial observation and maintains, or even improves, performance under full observation.